By the end of the Cold War, the Soviet enrichment facility UEKhk (now Novouralsk) reached a capacity of 12 million SWU/year. Now the facility has a production capacity of 10 million SWU/year. The facility is said to enrich to a maximum of 30% U-235 for research and fast breeder reactors. How long would it take for the facility to convert enough of the 30% enriched U to construct a nuclear weapon with 95% enrichment? State any assumptions.
Answer #1

Capacity = 12 million SWU/ year

Production: 10 million SWU/ year

Currently, the U-235 is 30% enriched

The next cycle (year) will enrich the U-235 further i.e. it will be 30%*(1+30%) = 39% enriched.

We need to find the number of cycles which will make the U-235, 95% enriched.

Let the number of cycles be n

i.e. 30%*(1+30%)*......*(1+30%) = 95% where the factor (1+30%) is repeated n times

i.e. 30% *(1.3)^n = 95%

i.e. 1.3^n = 95%/ 30% =3.167

i.e. n = ln 3.167/ ln (1.3)

i.e. n =1.152785/ 0.262364

i.e. n = 4.4 cycles